Please help me solve this equation i need to find the value of x of a rectangle l= x+3 and w = 2x-6 the perimeter = 26

Answer by macston(5194)

.
L=length=x+3; W=width=2x-6; P=perimeter=26
.
P=2(L+W)
26=2(L+W)
13=L+W
13=(x+3)+(2x-6)
13=3x-3
16=3x
16/3=x
ANSWER: x=16/3